Question 4:

The The advertised weight on a can of soup is 20 ounces. The actual weight in the cans follows a uniform distribution and varies between 19.3 and 20.3 ounces.

a.

What is the probability that a can of soup will have between 19.4 and 20.3 ounces?

b.

What is the mean weight of a can of soup?

c.

What is the standard deviation of the weight?
Answer #1

Solution :

Given that,

a = 19.3

b = 20.3

a) P(c < x < d) = (d - c) / (b - a)

P(19.4 < x < 20.3) = (20.3 - 19.4) / (20.3 - 19.3)

P(19.4 < x < 20.3) = 0.9

b) mean = (a + b) / 2

mean = (19.3 + 20.3) / 2

mean = 19.8

c) standard deviation =  (b - a) / 12

standard deviation =  (20.3 - 19.3) / 12

standard deviation =0.2887